OpenStudy (anonymous):

I need to find 2 points on a linear equation y=5 over 4x - 5

OpenStudy (anonymous):

cant get you

OpenStudy (anonymous):

why not

OpenStudy (anonymous):

\[y = (5/4)x-5\]

OpenStudy (anonymous):

I'm having a hard time understanding these equations

OpenStudy (anonymous):

choose any two numbers, lievy, and tell me what those numbers are

OpenStudy (anonymous):

0 and 1

OpenStudy (anonymous):

its y=5/(4x-5) and so plug in a value for x and get the corresponding value for y....u'll hav ur two points.

OpenStudy (anonymous):

okay, we substitute x =0 and get y =-5 so(0.-5) is one point on the given line similarly we substitute x =1 and end up with y = -15/4 so (1,-15/4) is another point on the given line

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you can have infinite number of points on that line. all you have to do is think of a number for x and get the corresponding value for y

OpenStudy (anonymous):

ok